# Reward-free World Models for Online Imitation Learning

### v1

@@ 1 Introduction
Motivated by these insights, we explore the application of world models in the context of online imitation learning without rewards, enabling IL agents to develop a deeper understanding of environmental dynamics and improve their performance in tasks characterized by high-dimensional observations and complex dynamics. In this work, we present a novel approach to online imitation learning that leverages the strengths of decoder-free world models, specifically designed for complex tasks involving high-dimensional observations, intricate dynamics, and vision-based inputs. In contrast to conventional latent world models, which rely on reward and Q-function estimation, our approach completely eliminates the need for explicit reward modeling. We propose a framework for reward-free world models that redefines the optimization process within the Q-policy space, addressing the instability associated with min-max optimization in the reward-policy space. By utilizing an inverse soft-Q learning objective for the critic network (Garg et al., 2021), our method derives rewards directly from Q-values and the policy, effectively rendering the world model reward-free. This novel formulation mitigates key challenges in IL, including out-of-distribution errors and bias accumulation.
@@ 2 Related Works
Recent works regarding IL leveraged deep neural architectures to achieve better performance. Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) formulated the reward learning as a min-max problem similar to GAN (Goodfellow et al., 2014). Inverse Soft Q-Learning (Garg et al., 2021) reformulated the learning objective of GAIL and integrated their findings into soft actor-critic (Haarnoja et al., 2018) and soft Q-learning agents for imitation learning. CFIL (Freund et al., 2023) introduced a coupled flow approach for reward generation and policy learning using expert demonstrations. ValueDICE (Kostrikov et al., 2019) proposed an off-policy imitation learning approach by transforming the distribution ratio estimation objective. Das et al. (2021) proposed a model-based inverse RL approach by predicting key points for imitation learning tasks. SQIL (Reddy et al., 2019) proposed an online imitation learning algorithm with soft Q functions. Diffusion Policy (Chi et al., 2023) is a recent offline IL method using a diffusion model for behavioral cloning. Implicit BC (Florence et al., 2022) discovers that treating supervised policy learning with an implicit model generally improves the empirical performance for robot learning tasks.
@@ 3 Preliminary
We model the decision-making process in the environment as a Markov Decision Process (MDP), which can be defined as a tuple $\langle\mathcal{S},\mathcal{A},p_{0},\mathcal{P},r,\gamma\rangle$. $\mathcal{S}$ and $\mathcal{A}$ represent state and action space. $p_{0}$ is the initial state distribution and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the transition probability. $r(s,a)\in\mathcal{R}$ is the reward function and $\mathcal{R}$ is the reward space. $\gamma\in(0,1)$ is the discount factor. We denote the expert state-action distribution as $\rho_{E}$ and the behavioral distribution as $\rho_{\pi}$. Similarly, we denote the expert policy as $\pi_{E}$ and the behavioral policy as $\pi$. $\Pi$ is the set of all stochastic stationary policies that sample an action $a\in\mathcal{A}$ given a state $s\in\mathcal{S}$. $\mathcal{Z}$ is the space for the latent representation of the original state observations, and $\mathcal{Q}$ is the space for all possible Q functions. $\mathcal{H}(\cdot)$ represents the entropy of a distribution.
@@
Inverse Reinforcement Learning (IRL) focuses on recovering a specific reward function $r(s,a)$ in the reward space $\mathcal{R}$ given a certain amount of expert samples using expert policy $\pi_{E}$. Maximum entropy IRL seeks to solve this problem by optimizing ${\text{max}}_{r\in\mathcal{R}}{\text{min}}_{\pi\in\Pi}\mathbb{E}_{\rho_{E}}[r(\mathbf{s},\mathbf{a})]-(\mathbb{E}_{\rho_{\pi}}[r(\mathbf{s},\mathbf{a})]+\mathcal{H}(\pi))$. GAIL (Ho & Ermon, 2016) generalized the objective into a form including an explicit reward mapping with a convex regularizer $\psi(r)$:
@@
which is the inverse soft-Q objective for critic learning. In the objective function, $p_{0}$ is the initial distribution of states. In this way, we can perform imitation learning by leveraging actor-critic architecture. The critic and policy can be learned by finding the saddle point in a joint optimization problem $Q^{*}=\text{argmax}_{Q\in\mathcal{Q}}\min_{\pi\in\Pi}\mathcal{J}(\pi,Q)$ and $\pi^{*}=\text{argmin}_{\pi\in\Pi}\max_{Q\in\mathcal{Q}}\mathcal{J}(\pi,Q)$. Garg et al. (2021) proved the uniqueness of the saddle point. For a fixed Q, the optimization for policy has a closed-form solution, which is the softmax policy:
@@
where $\phi$ is a concave function. Specifically, if we leverage $\chi^{2}$ regularization, we will have $\phi(x)=x-\frac{1}{4\alpha}x^{2}$, and $\alpha$ is a scalar coefficient to ensure stable training and prevent overfitting.
@@ 4 Methodology
In conventional latent world models, an explicit reward function is typically used to map latent representations and actions to observed reward data. However, in imitation learning, the model is limited to learning from a finite set of expert state-action demonstrations, making it difficult for traditional world models to directly perform imitation learning. To address this limitation, we propose a reward-free world model that learns solely from expert demonstrations and environment interactions, without requiring explicit reward data. Additionally, our model can decode dense rewards for the environment, allowing it to learn a reward function through state-only interactions and a limited set of expert demonstrations. This allows our model to effectively address inverse reinforcement learning (IRL) tasks.
@@ 4.1 Learning Process of a Reward-free World Model
World models used in reinforcement learning settings often contain a reward model $R(\mathbf{z},\mathbf{a})$ that requires supervised learning using explicit reward signals from the online environment interactions or the offline data. However, if our learning objective is able to form a bijection between Q space $\mathcal{Q}$ and reward space $\mathcal{R}$, it would be natural to decode the reward from the Q value instead of learning another separate mapping for the reward, which also enables the world model to perform imitation learning with expert demonstrations without explicit reward signal. An overview of our proposed method is shown in Figure 1. The detailed training algorithm is shown in Algorithm 2.
@@
Compared to Eq.6, the key difference is the second term of the objective, which computes the original value difference $\mathbb{E}_{(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}^{\prime}_{t})\sim\mathcal{B}_{\pi}}[V^{\pi}(\mathbf{z}_{t})-\gamma V^{\pi}(\mathbf{z}^{\prime}_{t})]$ using only the representation of the initial state $\mathbf{s}_{0}$. This reformulation, derived in Theorem 1 (Appendix F.1), yields more stable Q estimation, as confirmed by the ablation in Appendix E.3. We also apply $\chi^{2}$ regularization, as noted in Garg et al. (2021). We leverage $\lambda\in(0,1]$ as a constant discounting weight over the horizon, guaranteeing the influence to be smaller for states and actions farther ahead. Note that $\lambda$ here differs from the environment discount factor $\gamma$. All value functions in the objective are computed from Q and policy network via $V^{\pi}(\mathbf{z})=\mathbb{E}_{\mathbf{a}\sim\pi(\cdot|\mathbf{z})}[Q(\mathbf{z},\mathbf{a})-\beta\log\pi(\mathbf{a}|\mathbf{z})]$, where $\beta$ is the entropy coefficient. We will further discuss the selection of $\beta$ in the policy learning part. Especially, $\bar{V}^{\pi}(h(\mathbf{s}^{\prime}))$ is the value function computed by the target Q network $\bar{Q}$. $\mathbf{z}$ is retrieved by rolling out dynamics model from the latent representation of the first state:
@@ 5 Experiments
We conduct experiments for locomotion and manipulation tasks to demonstrate the effectiveness of our approach. We choose to leverage the online version of IQ-Learn+SAC (referred to as IQL+SAC in the experiment plots) (Garg et al., 2021) and CFIL+SAC (Freund et al., 2023) as our baselines for comparison studies. The empirical results regarding state-based and visual experiments are shown in Section 5.1. We experiment on the reward recovery capability of our IQ-MPC model, for which we reveal the results in Section 5.2. We conduct ablation studies for our model, which are discussed in Section 5.3. The details of the environments and tasks can be found in Appendix D.
Figure 2: Locomotion Results Our method demonstrates much stabler performance near expert level compared to baseline methods. In the plots, blue lines refers to the online version of IQ-L+SAC (Garg et al., 2021), orange lines refers to the CFIL+SAC method (Freund et al., 2023) and red lines refers to our IQ-MPC model. The dotted green lines are the mean episode reward for the expert trajectories used during training.
@@ 5.1.1 State-based Experiments
We benchmark our algorithm on DMControl (Tunyasuvunakool et al., 2020), evaluating tasks in both low- and high-dimensional environments. Our method outperforms baselines in performance and training stability. We use 100 expert trajectories for low-dimensional tasks (Hopper, Walker, Quadruped, Cheetah), 500 for Humanoid, and 1000 for Dog (both high-dimensional). Each trajectory contains 500 steps, sampled using trained TD-MPC2 world models (Hansen et al., 2023). Performance is averaged over 3 seeds per task. The results are demonstrated in Figure 2. We also conducted high-dimensional experiments on various tasks in the Dog environment, with results provided in Appendix E.1.
@@
Method IQL+SAC CFIL+SAC IQ-MPC(Ours)
Key Turn 0.72$\pm$0.04 0.65$\pm$0.08 0.87$\pm$0.03
Object Hold 0.00 $\pm$ 0.00 0.01$\pm$0.01 0.96$\pm$0.03
Pen Twirl 0.00 $\pm$ 0.00 0.00$\pm$0.00 0.73$\pm$0.05
@@
We consider manipulation tasks with a dexterous hand from MyoSuite (Caggiano et al., 2022) to show the capability and robustness of our IQ-MPC model in high-dimensional and complex dynamics scenarios. We leverage 100 expert trajectories with 100 steps sampled from trained TD-MPC2 for each task. We evaluate the episode reward and success rate of our model along with IQ-Learn+SAC and CFIL+SAC. We show superior empirical performance in three different tasks, including object holding, pen twirling, and key turning. Regarding the results for episode reward, we refer to Figure 3. Table 1 shows the success rate results. We take the mean for 3 seeds regarding the performance for each task. We have conducted additional experiments on ManiSkill2 (Gu et al., 2023), for which we refer to Appendix E.2.
Figure 4: Reward Recovery. The IQ-MPC model successfully recovers rewards in the inverse RL setting. The recovered rewards exhibit a positive correlation with ground truth rewards. This experiment is conducted on the Cheetah Run task using state-based observations from DMControl.
5.1.2 Visual Experiments
5.2 Reward Recovery

### v2

@@ 1 Introduction
Motivated by these insights, we explore the application of world models in the context of online imitation learning without rewards, enabling IL agents to develop a deeper understanding of environmental dynamics and improve their performance in tasks characterized by high-dimensional observations and complex dynamics. In this work, we present a novel approach to online imitation learning that leverages the strengths of decoder-free world models, specifically designed for complex tasks involving high-dimensional observations, intricate dynamics, and vision-based inputs. In contrast to conventional latent world models, which rely on reward and Q-function estimation, our approach completely eliminates the need for explicit reward modeling. We propose a framework for reward-free world models that redefines the optimization process within the Q-policy space, addressing the instability associated with min-max optimization in the reward-policy space. By utilizing an inverse soft-Q learning objective for the critic network (Garg et al., 2021), our method derives rewards directly from Q-values and the policy, effectively rendering the world model reward-free. Moreover, by performing imitation learning online, our model addresses key challenges in IL, such as out-of-distribution errors and bias accumulation.
@@ 2 Related Works
Recent works regarding IL leveraged deep neural architectures to achieve better performance. Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) formulated the reward learning as a min-max problem similar to GAN (Goodfellow et al., 2014). Model-based Adversarial Imitation Learning (MAIL) (Baram et al., 2016) extended the GAIL approach to incorporate a forward model trained by data-driven methodology. Inverse Soft Q-Learning (Garg et al., 2021) reformulated the learning objective of GAIL and integrated their findings into soft actor-critic (Haarnoja et al., 2018) and soft Q-learning agents for imitation learning. CFIL (Freund et al., 2023) introduced a coupled flow approach for reward generation and policy learning using expert demonstrations. ValueDICE (Kostrikov et al., 2019) proposed an off-policy imitation learning approach by transforming the distribution ratio estimation objective. Das et al. (2021) proposed a model-based inverse RL approach by predicting key points for imitation learning tasks. SQIL (Reddy et al., 2019) proposed an online imitation learning algorithm with soft Q functions. Diffusion Policy (Chi et al., 2023) is a recent offline IL method using a diffusion model for behavioral cloning. Implicit BC (Florence et al., 2022) discovers that treating supervised policy learning with an implicit model generally improves the empirical performance for robot learning tasks. Hybrid inverse reinforcement learning (Ren et al., 2024) proposed a new methodology leveraging a mixture of online and expert demonstrations for agent training, achieving robust performance in environments with stochasticity. Prior works (Englert et al., 2013; Hu et al., 2022; Igl et al., 2022) explored the potentials of model-based imitation learning on real-world robotics control and autonomous driving. EfficientImitate (Yin et al., 2022) combined EfficientZero (Ye et al., 2021) with adversarial imitation learning, achieving excellent results in DMControl (Tassa et al., 2018) imitation learning tasks. V-MAIL (Rafailov et al., 2021) introduced a model-based approach for imitation learning using variational models. CMIL (Kolev et al., 2024) proposed an imitation learning approach with conservative world models for image-based manipulation tasks. Ditto (DeMoss et al., 2023) developed an offline imitation learning approach with Dreamer V2 (Hafner et al., 2020) and adversarial imitation learning. DMIL (Zhang et al., 2023) utilized a discriminator to simultaneously evaluate both the accuracy of the dynamics and the suboptimality of model rollout data relative to real expert demonstrations in the context of offline imitation learning.
@@ 3 Preliminary
We model the decision-making process in the environment as a Markov Decision Process (MDP), which can be defined as a tuple $\langle\mathcal{S},\mathcal{A},p_{0},\mathcal{P},r,\gamma\rangle$. $\mathcal{S}$ and $\mathcal{A}$ represent state and action space. $p_{0}$ is the initial state distribution and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the transition probability. $r(\mathbf{s},\mathbf{a})\in\mathcal{R}$ is the reward function and $\mathcal{R}$ is the reward space. $\gamma\in(0,1)$ is the discount factor. We denote the expert state-action distribution as $\rho_{E}$ and the behavioral distribution as $\rho_{\pi}$. Similarly, we denote the expert policy as $\pi_{E}$ and the behavioral policy as $\pi$. $\Pi$ is the set of all stochastic stationary policies that sample an action $\mathbf{a}\in\mathcal{A}$ given a state $\mathbf{s}\in\mathcal{S}$. $\mathcal{Z}$ is the space for the latent representation of the original state observations, and $\mathcal{Q}$ is the space for all possible Q functions. $\mathcal{H}(\cdot)$ represents the entropy of a distribution.
@@
Inverse Reinforcement Learning (IRL) focuses on recovering a specific reward function $r(s,a)$ in the reward space $\mathcal{R}$ given a certain amount of expert samples using expert policy $\pi_{E}$. Maximum entropy IRL (Ziebart et al., 2008) seeks to solve this problem by optimizing ${\text{max}}_{r\in\mathcal{R}}{\text{min}}_{\pi\in\Pi}\mathbb{E}_{\rho_{E}}[r(\mathbf{s},\mathbf{a})]-(\mathbb{E}_{\rho_{\pi}}[r(\mathbf{s},\mathbf{a})]+\mathcal{H}(\pi))$. GAIL (Ho & Ermon, 2016) generalized the objective into a form including an explicit reward mapping with a convex regularizer $\psi(r)$:
@@
which is the inverse soft-Q objective for critic learning. In this way, we can perform imitation learning by leveraging actor-critic architecture. The critic and policy can be learned by finding the saddle point in a joint optimization problem $Q^{*}=\text{argmax}_{Q\in\mathcal{Q}}\min_{\pi\in\Pi}\mathcal{J}(\pi,Q)$ and $\pi^{*}=\text{argmin}_{\pi\in\Pi}\max_{Q\in\mathcal{Q}}\mathcal{J}(\pi,Q)$. Garg et al. (2021) proved the uniqueness of the saddle point. For a fixed Q, the optimization for policy has a closed-form solution, which is the softmax policy:
@@
where $\phi$ is a concave function. Specifically, if we leverage $\chi^{2}$ regularization, we will have $\phi(x)=x-\frac{1}{4\alpha}x^{2}$. The scalar coefficient $\alpha$ controls the strength of $\chi^{2}$ regularization in the inverse soft-Q objective. Intuitively, the additonal regularization term penalizes the magnitude of the estimated reward. Prior work (Al-Hafez et al., 2023) interpreted the objective with this regularizer as minimizing the squared Bellman error, establishing a connection between inverse soft-Q learning and SQIL (Reddy et al., 2019). A detailed empirical analysis on hyperparameter $\alpha$ is shown in Appendix E.3.
@@ 4 Methodology
In reinforcement learning, world models typically regress explicit reward signals provided by the environment. In imitation learning, prior approaches (Kolev et al., 2024; DeMoss et al., 2023) aim to train a reward model through adversarial objectives alongside a separate critic network trained on temporal difference objectives. In contrast, we eliminate the need for a separate reward model by retrieving rewards directly from the learned critic. To this end, we propose a reward-free world model that learns exclusively from reward-free expert demonstrations and environment interactions, without training a dedicated reward model. Furthermore, since our model can decode dense rewards from the critic, it can solve inverse reinforcement learning tasks using reward-free interactions and a limited set of expert demonstrations that include only states and actions.
@@ 4.1 Learning Process of a Reward-free World Model
World models used in reinforcement learning settings often contain a reward model $R(\mathbf{z},\mathbf{a})$ that requires supervised learning using explicit reward signals from the online environment interactions or the offline data. However, if our learning objective is able to form a bijection between Q space $\mathcal{Q}$ and reward space $\mathcal{R}$, it would be natural to decode the reward from the Q value instead of learning another separate mapping for the reward, which also enables the world model to perform imitation learning with expert demonstrations without explicit reward signal. An overview of our proposed method is shown in Figure 1. The detailed training algorithm is shown in Algorithm 2. We also provide a theoretical analysis of our training objective, as detailed in Section 4.2 Appendix G.3.
@@
Compared to Eq.6, the key difference is the second term of the objective, which computes the original value difference $\mathbb{E}_{(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}^{\prime}_{t})\sim\mathcal{B}_{\pi}}[V^{\pi}(\mathbf{z}_{t})-\gamma V^{\pi}(\mathbf{z}^{\prime}_{t})]$ using only the representation of the initial state $\mathbf{s}_{0}$. This reformulation, derived in Lemma 2 (Appendix G.1), yields more stable Q estimation, as confirmed by the ablation in Appendix E.3. We also apply $\chi^{2}$ regularization, as noted in Garg et al. (2021). We leverage $\lambda\in(0,1]$ as a constant discounting weight over the horizon, guaranteeing the influence to be smaller for states and actions farther ahead. Note that $\lambda$ here differs from the environment discount factor $\gamma$. All value functions in the objective are computed from Q and policy network via $V^{\pi}(\mathbf{z})=\mathbb{E}_{\mathbf{a}\sim\pi(\cdot|\mathbf{z})}[Q(\mathbf{z},\mathbf{a})-\beta\log\pi(\mathbf{a}|\mathbf{z})]$, where $\beta$ is the entropy coefficient. We will further discuss the selection of $\beta$ in the policy learning part. Especially, $\bar{V}^{\pi}(h(\mathbf{s}^{\prime}))$ is the value function computed by the target Q network $\bar{Q}$. $\mathbf{z}$ is retrieved by rolling out dynamics model from the latent representation of the first state:
@@
4.2 Theoretical Analysis on the Learning Objective
In this section, we demonstrate theoretically that the learning objectives of IQ-MPC effectively minimize the value difference between the current policy and the expert, ensuring that Q-value estimation can follow as the latent dynamics model learns. We begin by utilizing the following lemma established in (Kolev et al., 2024):
Lemma 1 (Bounded Suboptimality).
Given an unknown latent MDP $\mathcal{M}$ and our learned latent MDP $\hat{\mathcal{M}}$ with transition probabilities $d$ and $\hat{d}$ in the latent state space $\mathcal{Z}$ and action space $\mathcal{A}$, and letting $R_{\max}$ denote the maximum reward of the unknown MDP, the difference between the expected return of the current policy, $\eta^{\pi}_{\mathcal{M}}$, and that of the expert policy, $\eta^{\pi_{E}}_{\mathcal{M}}$, is bounded by:
$|\eta^{\pi_{E}}_{\mathcal{M}}-\eta^{\pi}_{\mathcal{M}}|\leq\underbrace{\frac{2R_{\max}}{1-\gamma}D_{TV}(\rho^{\pi}_{\hat{\mathcal{M}}},\rho^{\pi_{E}}_{{\mathcal{M}}})}_{\text{T1}}+\underbrace{\frac{\gamma R_{\max}}{(1-\gamma)^{2}}\mathbb{E}_{\rho^{\pi}_{\hat{\mathcal{M}}}}\Big{[}D_{TV}(d(\mathbf{z}^{\prime}|\mathbf{z},\mathbf{a}),\hat{d}(\mathbf{z}^{\prime}|\mathbf{z},\mathbf{a}))\Big{]}}_{\text{T2}}$
Our critic and policy objectives can be interpreted as a min-max optimization of Eq.4. Moreover, this approach can be viewed as minimizing a statistical distance with an entropy term, corresponding to Eq.2. Thus, on one hand, our critic and policy objectives effectively minimize T1 in the bound provided in Lemma 1. On the other hand, optimizing our consistency loss in Eq.11 is approximately minimizing the second term T2 in the bound. Our training objective ensures that as the dynamics model learns, it simultaneously minimizes the upper bound of the deviation between the expected return for current policy $\pi$ and the expert expected return. A more detailed analysis on T2 is given in Appendix G.3.
@@ 5 Experiments
We conduct experiments for locomotion and manipulation tasks to demonstrate the effectiveness of our approach. We choose to leverage the online version of IQ-Learn+SAC (referred to as IQL+SAC in the experiment plots) (Garg et al., 2021), CFIL+SAC (Freund et al., 2023), and HyPE (Ren et al., 2024) as our baselines for comparison studies. The results presented below for our IQ-MPC model are obtained through planning. We provide an analysis of the computational overhead of our model in Appendix F. The empirical results regarding state-based experiments are shown in Section 5.1. For visual results, we refer to Appendix E.4. We experiment on the reward recovery capability of our IQ-MPC model, for which we reveal the results in Appendix E.6. We conduct ablation studies for our model, which are discussed in Section 5.2. The details of the environments and tasks can be found in Appendix D. We also analyze the training time and the robustness of our model under noisy environment dynamics. The corresponding results are presented in Appendix F and E.5.
Figure 2: Locomotion Results Our method demonstrates much stabler performance near expert level compared to baseline methods. In the plots, blue lines refers to the online version of IQ-L+SAC (Garg et al., 2021), orange lines refers to the HyPE method (Ren et al., 2024), purple lines refers to the CFIL+SAC (Freund et al., 2023) baseline and red lines refers to our IQ-MPC model. The dotted green lines are the mean episode reward for the expert trajectories used during training.
@@ 5.1.1 State-based Experiments
We benchmark our algorithm on DMControl (Tunyasuvunakool et al., 2020), evaluating tasks in both low- and high-dimensional environments. Our method outperforms baselines in performance and training stability. We use 100 expert trajectories for low-dimensional tasks (Hopper, Walker, Quadruped, Cheetah), 500 for Humanoid, and 1000 for Dog (both high-dimensional). Each trajectory contains 500 steps, sampled using trained TD-MPC2 world models (Hansen et al., 2023). Performance is averaged over 3 seeds per task. The results are demonstrated in Figure 2. Our method is comparable to HyPE in the Quadruped Run and Cheetah Run tasks, while outperforming all other baselines in the remaining tasks. We also conducted high-dimensional experiments on various tasks in the Dog environment, with results provided in Appendix E.1.
@@
Method IQL+SAC CFIL+SAC HyPE IQ-MPC(Ours)
Key Turn 0.72$\pm$0.04 0.65$\pm$0.08 0.55 $\pm$ 0.09 0.87$\pm$0.03
Object Hold 0.00 $\pm$ 0.00 0.01$\pm$0.01 0.13 $\pm$ 0.10 0.96$\pm$0.03
Pen Twirl 0.00 $\pm$ 0.00 $0 0$$\pm$ 0.00 0.73$\pm$0.05
@@
We consider manipulation tasks with a dexterous hand from MyoSuite (Caggiano et al., 2022) to show the capability and robustness of our IQ-MPC model in high-dimensional and complex dynamics scenarios. We leverage 100 expert trajectories with 100 steps sampled from trained TD-MPC2 for each task. We evaluate the episode reward and success rate of our model along with IQ-Learn+SAC, HyPE, and CFIL+SAC. We show superior empirical performance in three different tasks, including object holding, pen twirling, and key turning. Regarding the results for episode reward, we refer to Figure 3. Table 1 shows the success rate results. We take the mean for 3 seeds regarding the performance for each task. We have conducted additional experiments on ManiSkill2 (Gu et al., 2023), for which we refer to Appendix E.2.